OpenStudy (anonymous):

Sorry For Reposting But The Answer Isnt Ellipse :(so please answer. Identify the conic section that is represented by the equation x2 3y2 + 6x + 2y - 5 = 0 Hyperbola Paraabola Ellipse Circle

OpenStudy (anonymous):

Rewrite as x^2 + 6x - [ 3y^2 - 2y ] = 5 Complete the squares: x^2 + 6x + 9 - [ y^2 - 2y/3 + 1/9 ] = 5 + 9 - 1/9 [ x + 3 ]^2 - [y - 1/3 ]^2 = 125/9 This is a hyperbola.

OpenStudy (anonymous):

THANK YOU :)

OpenStudy (anonymous):

your welcome :)